What is the length of the wooden strip required to frame a photograph of length and breadth 32 cm and 21 cm respectively?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Required length of wooden strip = Perimeter of photograph
= 2 (Length + Breadth) = 2 (32 + 21)
= 2 (53)
= 2 × 53
= 106 cm
∴Required length of the wooden strip is 106 cm

Thanks